OpenStudy (anonymous):

A lawn mower manufacturer incurs a total of $34,816 in overhead costs and $388 per lawn mower in production costs. How many lawn mowers were manufactured if the average cost of production is $660? The number of lawn mowers manufactured was_______?

OpenStudy (anonymous):

660

OpenStudy (anonymous):

@baru

OpenStudy (baru):

aagh wait, lag

OpenStudy (baru):

let \(n\) be the number of lawn movers total cost= overhead + (production cost per lawn mover \(\times n\)) total cost=34816 + (\(388 n\)) this is equation 1

OpenStudy (baru):

total cost= average cost \(\times n\) total cost= 660n this is equation 2

OpenStudy (baru):

equation 1 and 2 are both equal to total cost 34816+388n= 660n

OpenStudy (anonymous):

Thank you for explaining how to solve it.

OpenStudy (baru):

sure :)